Calculate the mass of 0.75moles of ammonium nitrate crystals (N=14, H=1, O=16)
1
Expert's answer
2020-05-06T13:57:49-0400

Solution.

"M(NH4NO3) = 14 + 4 + 14 + 48 = 50 \ g/mol"

"m(NH4NO3) = n(NH4NO3) \times M(NH4NO3)"

m(NH4NO3) = 37.5 g

Answer:

m(NH4NO3) = 37.5 g